OpenStudy (anonymous):

Find the solutions of the equation that are in the interval [0, 2pi) csc5 theta = 16 csc theta

OpenStudy (anonymous):

\[\csc(5\theta)=16\csc(\theta)\] or is it raised to the power of five?

OpenStudy (anonymous):

\[\csc^5(\theta)=16\csc(\theta)\]?